Ann walks 5 km/hr and betty walks 4 km/hr. If they start walking toward each other from a distance of 36 km, how many hours will it take for them to meet.
Answer by mananth(16946)

Ann walks 5 km/hr and betty walks 4 km/hr. If they start walking toward each other from a distance of 36 km, how many hours will it take for them to meet.
let them meet at a distance x from Ann
The distance from betty when they meet = 36-x
.
time taken to walk to the point of meeting by Ann = x/5
time taken by betty to walk to the point of meeting = 36-x/ 4
These times are same
x/5 = 36-x/4
4x=5(36-x)
4x=180-5x
9x = 180
x= 20 miles.
This is the distance traveled by Ann at 5 mph
So they meet after 20 / 5 = 4 hours